OpenStudy (anonymous):

What is the solution to this system of equations? 3x - 5y=-16 2x + 5y=31 A.(1/3, 1) B.(5,3) C.(8, 2) D.(3 ,5)

OpenStudy (anonymous):

alright for this elimination easiest way is to add. 3x-5y=-16 +2x+5y=31 The result should be 5x=15 and x=3 plug into any original equation 2(3)+5y=31= 6+5y=31 5y=25 y=5

OpenStudy (anonymous):

answer (3,5)

OpenStudy (anonymous):

Oh, hah thank you

OpenStudy (anonymous):

The reason we add in this case is becuz we have to eliminate something right? in this case if we add we would eliminate 5 and -5

OpenStudy (anonymous):

oh..

OpenStudy (anonymous):

it will get a little complicated later on when you get to different number where you have to multiply to get...you know the elimination process

OpenStudy (anonymous):

Are you already out of highschool?

OpenStudy (anonymous):

no are u

OpenStudy (anonymous):

no... I failed 9th grade so i'm taking summer classes so i can go to 10th.. i'm supposed to be in 11th..